# Twists of Superconformal Algebras

### v2

@@ 5.3.2. Relationship with Donaldson–Witten supercharges and real twists
Consider now the image $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$ of $[\mathcal{Q},-]$ inside the even part of the complexified superconformal algebra. It is isomorphic, as a Lie algebra, to $\mathfrak{sl}(2,\mathbb{C})\ltimes(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, and we are free to view the four dimensional abelian factor as the complexified translations of an affine slice, so let us denote this subalgebra by $T_{\mathbb{C}}$. (There is another copy of $\mathbb{C}^{4}$, orthogonal to this image, which we can then view as the special conformal transformations for this affine slice.)
Even better, as $\mathcal{Q}$ comes from a real form, it is a fixed point of the corresponding involution (i.e., of the appropriate version of complex conjugation). The bracket is equivariant for this involution, so the image of $[\mathcal{Q},-]$ restricted to the odd part of the real form gives us a 4-dimensional real subspace $T_{\mathbb{R}}$ of $T_{\mathbb{C}}$. We define an affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$ as the inclusion of the open orbit of the $T_{\mathbb{R}}$-action on the conformal compactification $C(\mathbb{R}^{p,4-p})$. ∎
@@
$\mathfrak{su}(2,2|2)=\{A\in\mathfrak{gl}(4|2,\mathbb{C})\,:\,A\mathbb{K}+(-1)^{|A|}\mathbb{K}A=0,\,\mathrm{str}(A)=0\}$
@@
with $Q\in\mathrm{Mat}_{2,2}(\mathbb{C})$ and $Q^{\dagger}$ its conjugate-transpose. Then
$[\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}KQ^{\dagger}QK&0\\ \hline\cr 0&QQ^{\dagger}\end{array}\right)$
by direct computation. Note that the diagonal entries of $QQ^{\dagger}$ are the norms of each row of $Q$, as a vector in $\mathbb{C}^{2}$ with the standard sesquilinear pairing. In other words, these diagonal entries are the norms as vectors in Euclidean $\mathbb{R}^{4}$ and hence must be nonnegative. We can then have $[\mathbb{Q},\mathbb{Q}]=0$ if and only if every row is zero, i.e., $Q=0$. ∎
@@
$[\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}BC&0\\ \hline\cr 0&CB\end{array}\right).$
@@ 6.1. Complex Nilpotence Variety
$[w\otimes Q,w^{\prime}\otimes Q^{\prime}]=(w\wedge w^{\prime})\omega(Q,Q^{\prime})+g(w,w^{\prime})(Q\cdot Q^{\prime})$
@@
be an arbitrary element of $(\mathfrak{sconf}(3|k,\mathbb{C}))_{1}$. Suppose this element squares to zero. If we first consider the $\mathfrak{so}(5,\mathbb{C})$ summand of $[\mathcal{Q},\mathcal{Q}]$, this means that
@@
Now, considering the $\mathfrak{so}(2k,\mathbb{C})$ summand of $[\mathcal{Q},\mathcal{Q}]$, we have
@@
$(\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(3,\mathbb{C})\oplus(\mathbb{C}^{3})^{2}\oplus\mathbb{C})\oplus\Pi(W_{k}\otimes(\mathbb{S}_{2})^{2}),$
@@ 6.2. Complex Group Orbits & Closed and Exact Transformations
acts on $\mathcal{N}^{(3)}_{k}$, and each complex orbit has a well-defined rank: that is, the rank of a square-zero supercharge $\mathcal{Q}\in W_{k}\otimes V_{4}$ viewed as a linear map $W_{k}\to V_{4}^{*}$. The image of the linear map associated to $\mathcal{Q}$ must be isotropic, so there will only be three complex orbits, associated to isotropic linear maps of rank 0, rank 1, and rank 2. (This classification is another example of the orbit classification of Duflo–Serganova [DufloSerganova, Theorem 4.2].)
For the rank $r$ complex orbit we will describe the kernel and the image of the operator $[\mathcal{Q},-]$ for a supercharge $\mathcal{Q}$ of rank $r$.
  • Rank 0: The zero supercharge always squares to zero. If $k<2$ this is the only point in $\mathcal{N}^{(3)}_{k}$. The kernel of $[\mathcal{Q},-]$ consists of the entirety of $\mathfrak{so}(5,\mathbb{C})\oplus\mathfrak{so}(k,\mathbb{C})$, and the image of $[\mathcal{Q},-]$ is just $\{0\}$.
@@
    The kernel $\mathfrak{z}_{\mathcal{Q}}$ of the operator $[\mathcal{Q},-]$ is isomorphic to the kernel of the map
@@
    embedded via the map $(A,X,Y)\mapsto(\mathrm{diag}(A,X),\mathrm{diag}(-A,Y))$. If $\lambda\mapsto A_{\lambda}$ under the natural isomorphism ×$\mathbb{C}^{\times}\cong\mathrm{SO}(2,\mathbb{C})$, certainly
@@
    The parabolic subalgebra $\mathfrak{p}_{k}\subseteq\mathfrak{so}(k,\mathbb{C})$ has nilradical of dimension $k-2$. To complete the argument we must show that the nilradical of $\mathfrak{p}_{5}\oplus\mathfrak{p}_{k}$ is contained in $\mathfrak{z}_{\mathcal{Q}}$. In other words, it is sufficient to show that the quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{l}$ is an abelian subgroup of the same dimension as the nilradical of $\mathfrak{p}$, which is exactly $(5-2)+(k-2)=k+1$. We can compute this directly. The quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{l}$ is spanned by matrices of the form
@@
    The image $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{z}_{\mathcal{Q}}$ of $[\mathcal{Q},-]$ is four-dimensional, spanned by a three-dimensional subspace of $\mathfrak{sp}(4,\mathbb{C})$ and a fourth generator contained in the diagonally embedded copy of $\mathfrak{so}(2,\mathbb{C})$ in the kernel described above.
@@
    Let us now consider the $\mathrm{SO}(k-2,\mathbb{C})$ action on the last $k-2$ coordinates. This action is transitive on the space of null vectors of the form $(0,0,z_{3},z_{4},\ldots,z_{k})$, so we may pass to a point in the same orbit where $z_{3}=1,z_{4}=i$ and $z_{j}=0$ for $j>4$.
@@ 6.3. Real Group Orbits
$\pi\colon\mathcal{O}_{r}\to\mathrm{SpGr}(4-r,4)$
where $\mathrm{SpGr}(r,4)$ is the Grassmannian of isotropic $r$-dimensional subspaces in $\mathbb{C}^{4}\cong T^{*}\mathbb{C}^{2}$, equipped with its standard holomorphic symplectic structure.
Now, we pursue the same approach as in Section 5.1. That is, we observe that the two factors $\mathrm{SO}(k,\mathbb{C})$ and $\mathrm{Sp}(4,\mathbb{C})$ of the even part of the superconformal group act only on the fiber and only on the base of the fibration $\pi$, respectively. That is, $\mathrm{SO}(k,\mathbb{C})$ acts on $\mathrm{GL}(r,\mathbb{C})$ by left multiplication, viewing $\mathrm{GL}(r,\mathbb{C})$ as the subgroup of $\mathrm{GL}(k,\mathbb{C})$ consisting of matrices with zeroes outside the upper-left $r\times r$ block. Meanwhile $\mathrm{Sp}(4,\mathbb{C})$ acts on the homogeneous space $\mathrm{SpGr}(4-r,4)\cong\mathrm{Sp}(4,\mathbb{C})/P$, again by left multiplication. Thus we may separate the question of orbits into the base and fiber directions. In the base we may use the analog of Theorem 5.5 for general semisimple groups.
Theorem 6.7 ([Wolf, Theorem 2.6]).
@@
Using the notation $\mathrm{Tr},\mathrm{red}$ from Section 4.1, the bracket is given by
$[Q\otimes w,Q^{\prime}\otimes w^{\prime}]=\big{(}F(Q\otimes Q^{\prime})\cdot\mathrm{Tr}(w\otimes w^{\prime}),\langle Q,Q^{\prime}\rangle\cdot\mathrm{red}(w\otimes w^{\prime})\big{)}.$
When we restrict $(\mathfrak{sconf}(5|2,\mathbb{C}))_{1}$ to a representation of $\mathfrak{so}(5,\mathbb{C})\subseteq\mathfrak{so}(7,\mathbb{C})$, we do indeed obtain two copies of $S_{(5)}\otimes W$, where $S_{(5)}$ is the four-dimensional (Dirac) spinor representation of $\mathfrak{so}(5,\mathbb{C})$. That is, two copies of the spinors in the 5d $\mathcal{N}=1$ supersymmetry algebra.
@@
In dimension 6 there exist $\mathcal{N}=(k,0)$ superconformal algebras for all positive integers $k$. We can identify superconformal algebras in dimension $6$ with the simple super Lie algebras $\mathfrak{osp}(8|2k)$ using the triality isomorphism to identify the defining representation of $\mathfrak{so}(8,\mathbb{C})$ with a (Weyl) spinor representation. Thus, writing $\mathfrak{sconf}(6|k,\mathbb{C})=\mathfrak{osp}(8|2k)$ for the $\mathcal{N}=k$ superconformal algebra, we can identify the even part with
@@
$[v\otimes w,v^{\prime}\otimes w^{\prime}]=(v\wedge v^{\prime}\omega(w,w^{\prime}),g(v,v^{\prime})F(w\otimes w^{\prime})),$
where $F\colon W_{2k}\otimes W_{2k}\to\mathfrak{sp}(2k,\mathbb{C})$ is the canonical projection.
@@
It is interesting to note that the $\mathcal{N}=(2,0)$ complex superconformal algebra in six dimensions and the $\mathcal{N}=k$ complex superconformal algebra in three dimensions are actually isomorphic; the roles of the conformal and R-symmetry transformations are interchanged. Both are isomorphic to the simple super Lie algebra $\mathfrak{osp}(8|4,\mathbb{C})$. In particular, in the $\mathcal{N}{=}\hskip 1.00006pt(2,0)$ case the classification of complex orbits for the even part of the complexified superconformal group is identical to the classification from Section 6.
@@ 8. Twisted Observables and Factorization Algebras
There are many ways to approach QFT mathematically, and here we will use a framework that captures the spacetime-dependence of operators and operator products via operads. In fact, we will use two approaches drawn from [CG1, CG2], freely referencing those books below. The first approach uses prefactorization algebras and is highly flexible, allowing one to work on arbitrary manifolds and with operators whose support is an arbitrary open subspace. A key result of [CG2] is that a Lagrangian field theory on the manifold $M$ naturally produces a prefactorization algebra on $M$, so we have a wealth of examples. (That book also examines the perturbative quantization of such theories in Euclidean signature and shows these also yield prefactorization algebras.) The second approach restricts to $M=\mathbb{R}^{n}$ and to open disks therein. It matches closely with standard manipulations in the physics literature and has the additional virtue that there are convenient comparison results letting one extract more familiar algebraic structures – like vertex algebras or associative algebras – efficiently. It may be more intuitive for most readers.
Both approaches let us make mathematical statements that are sharp versions of the discussion in Section 3. Our discussion here amounts, primarily, to mimicking ideas and results from [CG1, ElliottSafronov], which examine chiral conformal and supersymmetric theories, respectively. The arguments carry over straightforwardly to superconformal theories, so we often state the result and merely indicate what needs to be changed.
In the final subsection we discuss results of Saberi and Williams [SaberiWilliams], who take a different approach to the Beem et al twists and exhibit a huge collection of examples, including Lagrangian 4d $\mathcal{N}=2$ superconformal field theories, that yield vertex algebras.

### v3

@@ 5.3.2. Relationship with Donaldson–Witten supercharges and real twists
Consider now the image $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$ of $\mathcal{Q},-]$ inside the even part of the complexified superconformal algebra. It is isomorphic, as a Lie algebra, to $\mathfrak{sl}(2,\mathbb{C})\ltimes(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, and we are free to view the four dimensional abelian factor as the complexified translations of an affine slice, so let us denote this subalgebra by $T_{\mathbb{C}}$. (There is another copy of $\mathbb{C}^{4}$, orthogonal to this image, which we can then view as the special conformal transformations for this affine slice.)
Even better, as $\mathcal{Q}$ comes from a real form, it is a fixed point of the corresponding involution (i.e., of the appropriate version of complex conjugation). The bracket is equivariant for this involution, so the image of $\mathcal{Q},-]$ restricted to the odd part of the real form gives us a 4-dimensional real subspace $T_{\mathbb{R}}$ of $T_{\mathbb{C}}$. We define an affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$ as the inclusion of the open orbit of the $T_{\mathbb{R}}$-action on the conformal compactification $C(\mathbb{R}^{p,4-p})$. ∎
@@
$\mathfrak{su}(2,2|2)=\{A\in\mathfrak{gl}(4|2,\mathbb{C})\,:\,A\mathbb{K}+(-1)^{|A|}\mathbb{K}A^{\dagger}=0,\,\mathrm{str}(A)=0\}$
@@
with $Q\in\mathrm{Mat}_{2,4}(\mathbb{C})$ and $Q^{\dagger}$ its conjugate-transpose. Then
$\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}KQ^{\dagger}QK&0\\ \hline\cr 0&QQ^{\dagger}\end{array}\right)$
by direct computation. Note that the diagonal entries of $QQ^{\dagger}$ are the norms of each row of $Q$, as a vector in $\mathbb{C}^{2}$ with the standard sesquilinear pairing. In other words, these diagonal entries are the norms as vectors in Euclidean $\mathbb{R}^{4}$ and hence must be nonnegative. We can then have $\mathbb{Q},\mathbb{Q}]=0$ if and only if every row is zero, i.e., $Q=0$. ∎
@@
$\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}BC&0\\ \hline\cr 0&CB\end{array}\right).$
@@ 6.1. Complex Nilpotence Variety
$w\otimes Q,w^{\prime}\otimes Q^{\prime}]=(w\wedge w^{\prime})\omega(Q,Q^{\prime})+g(w,w^{\prime})(Q\cdot Q^{\prime})$
@@
be an arbitrary element of $(\mathfrak{sconf}(3|k,\mathbb{C}))_{1}$. Suppose this element squares to zero. If we first consider the $\mathfrak{so}(5,\mathbb{C})$ summand of $\mathcal{Q},\mathcal{Q}]$, this means that
@@
Now, considering the $\mathfrak{so}(2k,\mathbb{C})$ summand of $\mathcal{Q},\mathcal{Q}]$, we have
@@
$(\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(3,\mathbb{C})\oplus(\mathbb{C}^{3})^{2}\oplus\mathbb{C})\oplus\Pi(W_{k}\otimes(\mathbb{S}_{2}\oplus\mathbb{S}_{2})),$
@@ 6.2. Complex Group Orbits & Closed and Exact Transformations
acts on $\mathcal{N}^{(3)}_{k}$, and each complex orbit has a well-defined rank: that is, the rank of a square-zero supercharge $\mathcal{Q}\in W_{k}\otimes V_{4}$ viewed as a linear map $W_{k}\to V_{4}^{*}$. The image of the linear map associated to $\mathcal{Q}$ must be isotropic, so there will only be three complex orbits, associated to isotropic linear maps of rank 0, rank 1, and rank 2. (This classification is another example of the orbit classification of Duflo–Serganova [20, Theorem 4.2].)
For the rank $r$ complex orbit we will describe the kernel and the image of the operator $\mathcal{Q},-]$ for a supercharge $\mathcal{Q}$ of rank $r$.
  • Rank 0: The zero supercharge always squares to zero. If $k<2$ this is the only point in $\mathcal{N}^{(3)}_{k}$. The kernel of $\mathcal{Q},-]$ consists of the entirety of $\mathfrak{so}(5,\mathbb{C})\oplus\mathfrak{so}(k,\mathbb{C})$, and the image of $\mathcal{Q},-]$ is just $\{0\}$.
@@
    The kernel $\mathfrak{z}_{\mathcal{Q}}$ of the operator $\mathcal{Q},-]$ is isomorphic to the kernel of the map
@@
    embedded via the map $(A,X,Y)\mapsto(\mathrm{diag}(A,X),\mathrm{diag}(-A,Y))$. If $\lambda\mapsto A_{\lambda}$ under the natural isomorphism $\mathbb{C}^{\times}\cong\mathrm{SO}(2,\mathbb{C})$, certainly
@@
    The parabolic subalgebra $\mathfrak{p}_{k}\subseteq\mathfrak{so}(k,\mathbb{C})$ has nilradical of dimension k2k-2. To complete the argument we must show that the nilradical of $\mathfrak{p}_{5}\oplus\mathfrak{p}_{k}$ is contained in $\mathfrak{z}_{\mathcal{Q}}$. In other words, it is sufficient to show that the quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{l}$ is an abelian subgroup of the same dimension as the nilradical of $\mathfrak{p}$, which is exactly (52)+(k2)=k+1(5-2)+(k-2)=k+1. We can compute this directly. The quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{l}$ is spanned by matrices of the form
@@
    The image $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{z}_{\mathcal{Q}}$ of $\mathcal{Q},-]$ is four-dimensional, spanned by a three-dimensional subspace of $\mathfrak{sp}(4,\mathbb{C})$ and a fourth generator contained in the diagonally embedded copy of $\mathfrak{so}(2,\mathbb{C})$ in the kernel described above.
@@
    Let us now consider the $\mathrm{SO}(k-2,\mathbb{C})$ action on the last k2k-2 coordinates. This action is transitive on the space of null vectors of the form $(0,0,z_{3},z_{4},\ldots,z_{k})$, so we may pass to a point in the same orbit where $z_{3}=1,z_{4}=i$ and $z_{j}=0$ for $j>4$.
@@ 6.3. Real Group Orbits
Let $\mathrm{SpGr}(r,4)$ denote the Grassmannian of isotropic $r$-dimensional subspaces in $\mathbb{C}^{4}\cong T^{*}\mathbb{C}^{2}$, equipped with its standard holomorphic symplectic structure.
@@
$\pi\colon\mathcal{O}_{r}\to\mathrm{SpGr}(r,4)$
that sends $\mathcal{Q}$, viewed as a rank $r$ linear map $W_{k}^{*}\to V_{4}$, to its image $\mathcal{Q}(W_{k}^{*})\subset V_{4}$.
Now, we pursue the same approach as in Section 5.1. That is, we observe that the two factors $\mathrm{SO}(k,\mathbb{C})$ and $\mathrm{Sp}(4,\mathbb{C})$ of the even part of the superconformal group act only on the fiber and only on the base of the fibration $\pi$, respectively. That is, $\mathrm{SO}(k,\mathbb{C})$ acts on $\mathrm{GL}(r,\mathbb{C})$ by left multiplication, viewing $\mathrm{GL}(r,\mathbb{C})$ as the subgroup of $\mathrm{GL}(k,\mathbb{C})$ consisting of matrices with zeroes outside the upper-left $r\times r$ block. Meanwhile $\mathrm{Sp}(4,\mathbb{C})$ acts on the homogeneous space $\mathrm{SpGr}(r,4)\cong\mathrm{Sp}(4,\mathbb{C})/P$, again by left multiplication. Thus we may separate the question of orbits into the base and fiber directions. In the base we may use the analog of Theorem 5.5 for general semisimple groups.
Theorem 6.7 ([46, Theorem 2.6]).
@@
We give a brief examination of the five- and six-dimensional settings.
@@
The bracket is given by
$Q\otimes w,Q^{\prime}\otimes w^{\prime}]=\big(F(Q\otimes Q^{\prime})\cdot\mathrm{Tr}(w\otimes w^{\prime}),\langle Q,Q^{\prime}\rangle\cdot\mathrm{red}(w\otimes w^{\prime})\big)$
where we use again the notation $\mathrm{Tr},\mathrm{red}$ from Section 4.1.
When we restrict $(\mathfrak{sconf}(5|2,\mathbb{C}))_{1}$ to a representation of $\mathfrak{so}(5,\mathbb{C})\subseteq\mathfrak{so}(7,\mathbb{C})$, we do indeed obtain two copies of $S_{(5)}\otimes W$, where $S_{(5)}$ is the four-dimensional (Dirac) spinor representation of $\mathfrak{so}(5,\mathbb{C})$ (that is, two copies of the spinors in the 5d $\mathcal{N}=1$ supersymmetry algebra).
@@
In dimension 6 there exist $\mathcal{N}=(k,0)$ superconformal algebras for all positive integers $k$. We can identify superconformal algebras in dimension $6$ with the simple super Lie algebras $\mathfrak{osp}(8|2k)$ using the triality isomorphism to view the defining representation of $\mathfrak{so}(8,\mathbb{C})$ as a (Weyl) spinor representation. Thus, writing $\mathfrak{sconf}(6|k,\mathbb{C})=\mathfrak{osp}(8|2k)$ for the $\mathcal{N}=(k,0)$ superconformal algebra, we can identify the even part with
@@
$v\otimes w,v^{\prime}\otimes w^{\prime}]=(\omega(w,w^{\prime})v\wedge v^{\prime},\,g(v,v^{\prime})F(w\otimes w^{\prime})),$
where $F\colon W_{2k}\otimes W_{2k}\to\mathfrak{sp}(2k,\mathbb{C})$ is the canonical projection, $\omega$ is the symplectic pairing on $W_{2k}$, and $g$ is the symmetric pairing on $S_{+}$.
We note that, because the relevant super Lie algebra is of the type $\mathfrak{osp}$, the techniques from Section 6 can be applied to this situation.
Example 7.2.
Let $k=1$, so $\mathfrak{sconf}(6|k,\mathbb{C})=\mathfrak{sconf}(6|1,\mathbb{C})\cong\mathfrak{osp}(8|2)$. In this case, let $\{w_{1},w_{2}\}$ denote a basis for $W_{2}$ such that $\omega(w_{1},w_{2})=1$. An arbitrary odd element then has the form
$Q=v_{1}\otimes w_{1}+v_{2}\otimes w_{2}$
with $v_{1},v_{2}\in V_{8}$. Then
$Q,Q]=(v_{1}\wedge v_{2},\,\sum_{i,j}g(v_{i},v_{j})F(w_{i}\otimes w_{j}))$
so the first component vanishes only if $v_{1}\wedge v_{2}=0$. In other words, we need $v_{2}=tv_{1}$ for some scalar $t$. That means the second component vanishes only if $g(v_{1},v_{1})=0$. Hence the complex nilpotence variety for the 6d $\mathcal{N}=(1,0)$ superconformal algebra is
$\{0\}\sqcup\mathcal{(}N^{(6)}_{(1,0)}\setminus\{0\})\times\mathbb{C}^{\times}$
with
$\mathcal{N}^{(6)}_{(1,0)}=\{v\in V_{8}\colon g(v,v)=0\}$
the null cone in $V_{8}$. In words, there is the zero point and for any nonzero $v$ in the nullcone, we can set $v_{1}=v$ and we pick a scalar $t$ so that $v_{2}=tv_{1}$. There is a single non-zero orbit under the action of the complex group $\mathrm{Spin}(6,\mathbb{C})$.
@@
It is interesting to note that the $\mathcal{N}=(2,0)$ complex superconformal algebra in six dimensions and the $\mathcal{N}=8=k$ complex superconformal algebra in three dimensions are actually isomorphic; the roles of the conformal and R-symmetry transformations are interchanged. Both are isomorphic to the simple super Lie algebra $\mathfrak{osp}(8|4,\mathbb{C})$. In particular, in the $\mathcal{N}{=}\hskip 1.00006pt(2,0)$ case the classification of complex orbits for the even part of the complexified superconformal group is identical to the classification from Section 6.
@@ 8. Twisted Observables and Factorization Algebras
There are many ways to approach QFT mathematically, and here we will use a framework that captures the spacetime-dependence of operators and operator products via operads. In fact, we will use two approaches drawn from [15, 16], freely referencing those books below. The first approach uses prefactorization algebras and is highly flexible, allowing one to work on arbitrary manifolds and with operators whose support is an arbitrary open subspace. A key result of [16] is that a Lagrangian field theory on the manifold $M$ naturally produces a prefactorization algebra on $M$, so we have a wealth of examples. (That book also examines the perturbative quantization of such theories in Euclidean signature and shows these also yield prefactorization algebras.) The second approach restricts to $M=\mathbb{R}^{n}$ and only considers operators supported on open disks and how they multiply. It matches closely with standard manipulations in the physics literature and has the additional virtue that there are convenient comparison results letting one extract more familiar algebraic structures – like vertex algebras or associative algebras – efficiently. It may be more intuitive for most readers.
Both approaches let us make mathematical statements that are sharp versions of the discussion in Section 3. Our discussion here amounts, primarily, to mimicking ideas and results from [15, 23], which examine chiral conformal and supersymmetric theories, respectively. The arguments carry over straightforwardly to superconformal theories, so we often state the result and merely indicate what needs to be changed.
In the final subsection we discuss results of Saberi and Williams [40], who take a different approach to the Beem et al twists and exhibit a huge collection of examples, including Lagrangian 4d $\mathcal{N}=2$ superconformal field theories, that yield vertex algebras.